Two cyclists, 90 mi apart, start riding toward each other at the same time. One cycles twice as fast as the other. If they meet 2 h later, at what average speed is each cyclist traveling?
Answer by mananth(16949)

distance 90 miles
speed of one cyclist = x
other speed = 2x
time they drive =2 hours
since they are travelling towards each other the effective speed = 2x+x=3x
Time = distance /speed
90/3x = 2
6x = 90
x = 15 mph speed of one cyclist.
speed of other will be 2x = 2*15 = 30 mph.
...
m.ananth@hotmail.ca